OpenStudy (anonymous):

PLEASE HELP, WILL GIVE MEDALS. PLEASE Quick Computing Company produces calculators. They have found that the cost, c(x), of making x calculators is a quadratic function in terms of x. The company also discovered that it costs $45 to produce 2 calculators, $143 to produce 4 calculators, and $869 to produce 10 calculators. Find the total cost of producing 1 calculator.

OpenStudy (matlee):

Hi again, do you have multiple choice or is it typing?

OpenStudy (anonymous):

No Multiple choice ... sadly

OpenStudy (matlee):

Can you tell me the options?

OpenStudy (anonymous):

95

OpenStudy (anonymous):

explanation:

OpenStudy (anonymous):

c(x) = 9x^2 - 5x + 19

OpenStudy (anonymous):

X=1 since we are solving for one claculator

OpenStudy (anonymous):

But how can one be more than how much it costs to make two?

OpenStudy (anonymous):

solve it out and 1 calculator = 95

OpenStudy (anonymous):

Matlee did you come up with the same answer?? Sorry this just has to be 100% right. I can't miss anymore questions :'(

OpenStudy (matlee):

Ok let me check

OpenStudy (anonymous):

I found where someone worked out the same problem just with different numbers if that helps anyone figure it out

OpenStudy (matlee):

Yes on yahoo answers right?

OpenStudy (anonymous):

Yes this is the one of yahoo answers. KEEP IN MIND: different numbers, and for my problem you are solving for 1 calculator Using the points (2, 45) (4, 143) and (10, 869), we can plug them into the following system of 3 equations using the y = ax^2 + bx + c format: 45 = a(2)^2 + b(2) + c 143 = a(4)^2 + b(4) + c 869 = a(10)^2 + b(10) + c which simplifies to: 45 = 4a + 2b + c 143 = 16a + 4b + c 869 = 100a + 10b + c Solving the system, we get a = 9, b = -5, and c = 19. Thus the equation is: c(x) = 9x^2 - 5x + 19 If you have a TI graphing calculator, you can also enter the points by pressing Stat -> Edit and enter (2, 45) (4, 143) and (10, 869) into it. Go back and calculate the QuadReg of the points from the Calc tab and it will give you the same answer. Now that we know the function that will produce the price of production for any number of calculators, plug in x = 7 and it will give you the price to produce 7 calculators. c(x) = 9x^2 - 5x + 19 ==> c(7) = 9(7)^2 - 5(7) + 19 ==> c(7) = 441 - 35 + 19 ==> c(7) = 425 Therefore, it costs $425 to produce 7 calculators.

OpenStudy (matlee):

So if you replace all the sevens with Ones you should get

OpenStudy (matlee):

c(x) = 9x^2 - 5x + 19 ==> c(1) = 9(1)^2 - 5(1) + 19

OpenStudy (matlee):

So 9^2= 81 and 5(1)=5

OpenStudy (matlee):

c(1)=81-5+19

OpenStudy (anonymous):

So you agree it's $95 ?

OpenStudy (matlee):

wait let me check this first

OpenStudy (anonymous):

Okay take your time. The only thing I have to say is, why is one costing more than two

OpenStudy (matlee):

No idea but also making 4 calculator is like 40% more than to make 2 calculators times two

OpenStudy (matlee):

yes i agree 95

OpenStudy (anonymous):

Okay I am about to enter it. Wish me luck! If it's wrong nonetheless you deserve a medal and all the thanks for taking the time to help me

OpenStudy (anonymous):

It was wrong! but thank you for the help

OpenStudy (matlee):

Im so sorry i have no idea how that is wrong

OpenStudy (matlee):

Your welcome